# Expert-Token Resonance MoE: Bidirectional Routing with Efficiency Affinity-Driven Active Selection

### v3

@@ Abstract
Mixture-of-Experts (MoE) architectures have emerged as a paradigm-shifting approach for large language models (LLMs), offering unprecedented computational efficiency. However, these architectures grapple with challenges of token distribution imbalance and expert homogenization, impeding optimal semantic generalization. We propose a novel expert routing framework that incorporates: (1) An efficient routing mechanism with lightweight computation. (2) An adaptive bidirectional selection mechanism leveraging resonance between experts and tokens. (3) A module that determines the lower bounds of expert capacity based on dynamic token distribution analysis, specifically designed to address drop-and-pad strategies. It is also integrated with orthogonal feature extraction module and an optimized loss function for expert localization. This framework effectively reduces expert homogeneity while enhancing the performance of the expert selection module. Additionally, we introduce a local expert strategy that simultaneously improves load balancing and reduces network communication overhead. It achieves a 40% reduction in token processed by each expert without compromising model convergence or efficacy. When coupled with communication optimizations, the training efficiency improvements of 5.4% to 46.6% can be observed. After supervised fine-tuning, it exhibits performance gains of 9.7% to 14.1% across GDAD, GPQA, and TeleQnA benchmarks.
@@ Introduction
Large language models (LLMs) have shown exceptional proficiency in understanding deep structures and complex semantic relationships within language (Zhao et al. 2023). As these models scale up, their capabilities in language generation and logical comprehension are enhanced, but this comes at the cost of significant computational, communication, and storage demands (Jiang et al. 2024b). To scale models efficiently without disproportionately increasing computational costs, researchers have incorporated the Mixture-of-Experts (MoE) architecture into LLMs (Lepikhin et al. 2020). The MoE framework integrates multiple experts within the model, each tasked with processing specific types of inputs (Fedus, Zoph, and Shazeer 2022). For a given input, only a subset of experts is activated, allowing for more efficient use of computational resources (Du et al. 2022). Recently, several LLMs employing MoE structures, such as DeepSeek-V3 (Liu et al. 2024a) and Mixtral (Jiang et al. 2024a), have demonstrated outstanding performance on various leaderboards.
Despite the efficiency benefits of MoE in scaling model sizes, it introduces several new challenges and drawbacks (Shazeer et al. 2017). The conventional MoE model’s convergence and the experts’ generalization capabilities are heavily dependent on the design of the routing strategy, which easily leads to an imbalanced ”winner-takes-all” phenomenon among experts. The imbalance between excessively ”developed” experts and those lacking adequate training may compromise or even nullify the intended functionality of routing strategies. Recent studies address these challenges from multiple perspectives (Li et al. 2023). StableMoE (Dai et al. 2022) proposes a two-stage training approach to address the issue of routing fluctuation. This method involves training the routing network independently from the backbone model and utilizing a frozen, distilled routing mechanism to allocate tokens. Dynamic-MoE (Huang et al. 2024a) designs a dynamic routing Mixture-of-Experts (MoE) policy that evaluates the sufficiency of current experts while reducing activated parameters by 90%. The characteristics of classical gated routing lead to experts being unable to learn features mastered by other experts. To address this, MoDE (Xie et al. 2024) proposes moderate distillation between experts to mitigate the generalization problems caused by narrow learning paths. DYNMoE (Guo et al. 2024) introduces a unique gated routing mechanism capable of adaptively determining the number of activated experts through trainable expert thresholds, even allowing for the addition or removal of experts.
In addition to the classical token choice scenario, previous researches also propose work utilizing expert choice (EC). Google Brain introduces the EC routing algorithm (Zhou et al. 2022), which assigns experts with predetermined buffer capacities to the Top-k tokens to ensure load balance. The Brainformer (Zhou et al. 2023) also adopts this routing strategy, constructing a trainable gating matrix to project the input feature space onto scores corresponding to each expert. Then, each token is routed to the Top-k experts. This strategy is proven highly effective in achieving expert load balancing and enhancing expert learning outcomes. Autonomy-of-Experts models (Lv et al. 2025) design a novel MoE paradigm in which experts autonomously select themselves to process inputs by aware of its own capacity to effectively process a token.
The design of routing strategy is crucial to the MoE structure, while not all tokens may be suitable for training (Riquelme et al. 2021). In addition to data preprocessing techniques such as dataset cleaning and deduplication, previous studies have also considered how to discard certain tokens within the model. Early work introduced the concept of expert capacity (Lepikhin et al. ), which refers to the maximum number of tokens each expert can process at once. Tokens exceeding this capacity are discarded. Expert capacity helps to ensure load balance among experts while facilitating All-to-All communication implementation. However, in situations where it is uncertain whether a token contributes to training, there is a risk of discarding class-discriminative samples, potentially compromising the model’s training outcomes. DeepSeek-V2 (Liu et al. 2024a) designs a device-limited routing mechanism to bound MoE-related communication cost. DeepSeek-V3 (Liu et al. 2024a) pioneers an auxiliary-loss-free strategy to minimizes the performance degradation. This approach minimizes the constraints on expert specialization imposed by knowledge hybridity and knowledge redundancy. XMoE (Yang et al. 2024) achieves more precise router by implementing a threshold-based approach. If a token reaches the specified threshold, it is processed exclusively by a single expert while being discarded by other experts within the Top-k selection. This method allows for more nuanced token selection and processing. LocMoE (Li et al. 2024) leverages orthogonal routing weights to prevent token homogenization across different expert networks and introduces the Grouped Average Pooling (GrAP) layer (Wang, Zhang, and Du 2023) for token feature extraction. Under these conditions, LocMoE also provides the theoretical proof for the lower bound of expert capacity.
In this paper, we propose expert-token resonance, a mechanism consisting of an expert-token bidirectional selection router and the adaptive expert capacity strategy. The primary contributions of this paper are as follows:
@@
    Affinity-based Efficient Expert Routing via GrAP. By leveraging cosine similarity between tokens and gating weights to define affinity scores, our router effectively guides experts to focus on distinct token segments, mitigating the expert homogenization problem. Meanwhile, the GrAP design reduces computational complexity by a factor of 1/2D to 1/D compared to traditional MLPs (D denotes the dimension of the intermediate hidden layer). This integrated approach demonstrates both improved routing effectiveness and substantial computational efficiency.
@@
    Expert-token Bidirectional Selection. By integrating the concepts of expert choice router (ECR) and token choice router (TCR), we propose the adaptive bidirectional selection mechanism. Contrast to conventional router, the bidirectional selection router allows MoE to enhance the training success rate while considering expert capacity constraints. Its effectiveness has been theoretically validated.
@@
    The Adaptive Expert Capacity Bound. Setting an adaptive affinity threshold allows the lower bound of expert capacity to be significantly reduced. As training iterations increase, the information density of token features grows, causing the expert capacity to initially decrease and then stabilize. Ultimately, the training efficiency of MoE can be greatly enhanced.
Expert-token resonance mechanism adopts the state-of-the-art MoE model Mixtral 8$\times$7B as the backbone, and utilizes MindSpeed-LLM, MindSpeed, and Megatron-LM (Shoeybi et al. 2019) libraries for training on Ascend NPU clusters. Ascend designs a new computing architecture for LLM training and inference scenarios (Liao et al. 2021), boasting powerful low-bit computing capabilities. Experiments conducted on clusters with 32, 64, and 256 NPUs indicate that our approach improves training efficiency by 5.4% to 46.6% compared to the baseline, and by 2.9% to 13.3% compared to LocMoE. Model performance is enhanced by 9.7% to 14.1% compared to the baseline, and by 1.7% to 4.1% compared to LocMoE.
The rest of this paper is structured as follows: Section Method presents the methods proposed in this paper, along with theoretical evidence. Section Experiments analyzes the experimental results of our approach regarding training efficiency and model performance. The final section summarizes the content of this paper and offers an outlook on future improvements.
@@
With GrAP as the layer of feature extraction, the formulation of $\bm{W}_{\text{aff}}$ is as followed:
@@
We consider our affinity-driven active selection routing as a hybrid of TCR (Clark et al. 2022; Zhou et al. 2022) and ECR. As the name suggested, TCR lets each token choose its top-scored experts, and ECR lets each expert choose its top-scored tokens. Specifically, we use the result of the expert-token affinity metrics as the affinity score between tokens and experts. In conventional TCR routing strategy, the tokens are simply route to their Top-1 expert. In our hybird TCR+ECR routing strategy, experts also select tokens for processing from assigned tokens according to affinity scores:
@@
and then the expert to choose its $-\ell$ tokens where $\ell$ is determined by a threshold of the sum of affinity scores:
@@
Feed-forward network (FFN) layers are commonly employed in expert networks, allowing each expert to learn independently as a separate neural network, thus preventing interference between samples. This mechanism leads to a severe load imbalance, as experts frequently selected in the early stages are more likely to be chosen in later stages. To mitigate this skewness in token allocation, the auxiliary loss (Shazeer et al. 2017) has been proposed. Building upon the auxiliary loss, our work introduces a loss bias term based on data locality, represented as $L_{\mathrm{loc}}=\mu\mathrm{KL}(D_{\mathrm{c}}||D_{\mathrm{l}})=-\mu\int{D_{\mathrm{c}}(x)\ln[\frac{D_{\mathrm{l}}(x)}{D_{\mathrm{c}}(x)}]\mathrm{d}x}$, i.e., the Kullback-Leibler (KL) divergence of the current distribution $D_{\mathrm{c}}(x)$ and the fully localized distribution $D_{\mathrm{l}}(x)$. This loss term serves as a soft constraint, encouraging tokens to be sent to experts residing on the same node, thereby mitigating the substantial overhead incurred by partial inter-node communication.
@@
Under the premises of orthogonal gating weights and a data distribution approaching uniformity, the previous studies demonstrate that the expert capacity is closely related to the angle between the gating weights and tokens. For large scale of the activation, the lower bound of expert capacity is proven to exist and is represented as $C_{\mathrm{min}}=\frac{1}{n}\exp\{d\delta_{\mathrm{max}}^{2}/(2-\delta_{\mathrm{max}}^{2})\}$.
The hybrid TCR+ECR bidirectional selection routing, introduced in the model structure, is exemplified in the figure. If the feature fragment corresponding to the $k$-th dimension of the gating weight for a particular token is more prominent, then that token will be routed to the $k$-th expert. If among all tokens routed to the $k$-th expert, there is a certain probability of the presence of class-discriminative tokens, then the capacity $C$ must be set to a larger value to ensure the inclusion of sufficient class-discriminative tokens. The router proposed in this paper is a hybrid of TCR and ECR modes. After determining the expert to which a token will be routed, scores are calculated for the tokens assigned to each expert, and a $-\ell$ selection is performed, where $\ell^{*}$ is determined by a threshold of the sum of scores. Subsequent theoretical analysis will demonstrate the effectiveness of this hybrid routing scheme.
Figure 3: (a) The average composition of computation, communication, overlap, and idle with different schemes and cluster sizes. (b) The perplexity during training iterations with different schemes.
Theoretical Explanation
To explain the motivation of our method, we show some theoretical insights in this section. Our theoretical analyis is bulit on Chowdhury et al. (2023), where they make the following data assumption:
@@
Each input $\bm{x}\in\mathcal{R}^{sd}$ with $s$ tokens is comprised of one class-discriminative pattern $\bm{o}_{1},\dots,\bm{o}_{n}\in\mathcal{R}^{d}$, with each decides the label in $[n]$, and $s-1$ class-irrelevant patterns $\bm{r}\sim\mathcal{N}$ for certain distribution $\mathcal{N}$. For example, $\bm{x}=(\bm{r}_{1},\bm{r}_{2},\bm{o}_{1},\bm{r}_{3},\dots,\bm{r}_{s-1})$ has label $1$, where $\bm{r}_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N},\forall i\in[s-1]$.
@@
Phase 1: Router training (Chowdhury et al. 2023, Lemma 4.1 and Assumption 4.4), which makes class-discriminative patterns all to the corresponding expert. This process ensures that each expert only receives the class-discriminative tokens related to the specific class.
Phase 2: Expert training (Chowdhury et al. 2023, Theorem 4.2 and Theorme 4.5), which makes each expert learn to predict the label based on its class-discriminative inputs from Phase 1. This process is designed to establish each expert’s ability to handle and solve problems.
Hence, the traning of an input in the current step is valid if the class-discriminative patterns is correctly dispatched. To quantitatively measure the difference between TCR and ECR, we define training success rate of input motivated by the training process of MoE.
@@
$i\sim\mathrm{Unif}([s]),\bm{x}_{i}\sim\mathrm{Unif}\left(\{\bm{o}_{1},\dots,\bm{o}_{n}\}\right).$
@@
$\mathcal{P}(\delta_{\bm{o}_{i},i}\geq\delta_{\bm{x}_{j},i},\forall j\in[s])=p_{i}\geq 1/n,\forall i\in[n].$
@@
Assumption 3 assumes the valid token is uniformly distributed in training samples due to the massive amounts of data nowadays. Assumption 4 assumes the invalid tokens can be uniformly dispatched to experts since the invalid tokens do not provide supervised signal to router and experts in training. We consider such uniform settings are common assumptions in theoretical analyis. Now we compute the training success rate of TCR and ECR.
@@
[…]
@@
We explain the benefit of swicthing TCR to ECR during training based on Theorem 5 and feature distrution during training.
At the beginning of training, the model seldom learn the task. Then the feature of class-irrelevant tokens is nearly isotropy, e.g., uniformly distrbute around the sphere (see Appendix), leading to $q_{i}=\Theta(1)$. The succed rate of TCR with the form $C/s$ is better than ECR with the form $e^{-s}$. Thus we should choose TCR with a large capacity $C=\Theta(s)$ to improve the success rate of training samples.
After training for some iterations, the experts can roughly distinguish the class-irrelevant and discriminative patterns, leading to $q_{i}\ll 1$ or $sq_{i}\leq C^{*}$ for some $C^{*}>0$ (see Appendix). Then ECR with success rate nearly $1$ is better than TCR with the form $C/s$ as long as $C\geq 2C^{*}$. Thus we should choose ECR with a small capacity $C=\Theta(1)$ to improve the success rate of training samples.
Indeed, we find that Chowdhury et al. (2023, the definition of $\ell^{*}$) consider the ECR setting and verify the benefit in sample complexity. They assume the maximum number of class-irrelevant patches that are close to class-discriminative patches are bounded, which has similar effect as $C^{*}$ in our scene.
@@
The training framework employs the Communication Over Computation (CoC) optimization technique to address performance bottlenecks in LLM training. During forward propagation in LLMs, the ColumnParallelLinear and RowParallelLinear components involve sequentially dependent computation (matrix multiplication) and communication (collective operations like AllReduce, AllGather, and ReduceScatter). These dependencies lead to inefficient serial execution. CoC decomposes these tasks into finer-grained subtasks and merges computation and communication into single kernels, such as MATMUL_ALL_REDUCE and MATMUL_REDUCE_SCATTER, utilizing MTE’s remote memory access capabilities. This approach allows for pipeline-style parallel execution and overlapping of computation and communication, significantly enhancing overall efficiency.
@@
This study employs the Mixtral 8×7B model, incorporating our proposed approach. The Mixtral model, comprising 46.7 billion parameters and utilizing Group Query Attention (GQA), features 32 sparse expert blocks with 8 experts in the MoE Feedforward layer, where each token engages the top 2 experts for processing. Given the prevalence of long-text corpora in our application scenarios, we extended the sequence length to 32,768 and implemented tailored parallel strategies for cluster scales of 32N, 64N, and 256N, encompassing tensor, pipeline, data, and expert parallelism, with a consistent global batch size of 128. For the three cluster scales of 32N, 64N, and 256N, the parallel strategies are set as follows: 32N - tensor parallel (TP=4) / pipeline parallel (PP=4) / data parallel (DP=2) / expert parallel (EP=2), 64N - TP=8 / PP=4 / DP=2 / EP=2, and 256N - TP=8 / PP=8 / DP=4 / EP=2. Other details of experimental setup including datasets, environment, and metrics, can be seen in Appendix.
@@
As detailed in Section Method, we consistently use Top-1 routing to ensure the routing implementation aligns with our theoretical framework. The Baseline model utilizes a limited expert capacity mode instead of the groupedGEMM scheme, which avoids token dropping, with the capacity factor set to 1.1. LocMoE considers data distribution uniformity and estimates expert capacity using a lower bound formula derived from its theoretical conclusions in the first batch, maintaining it as a constant during subsequent training. Our approach (abbreviate to ”LocMoE+” in figures) fixes the range of score sums, processes hidden states, and calculates current expert capacity. The subsequent analysis addresses the training time, convergence, and memory usage efficiency of these schemes on multiple sizes of Ascend clusters.
@@
Figure 3a illustrates the time consumption of these methods during the first 1000 iterations of training. Due to initialization and some unstable factors, time consumption is recorded starting from the 5th iteration. The Baseline model’s time consumption is relatively stable. As iterations increase, LocMoE’s time consumption slightly decreases, particularly in 32N and 64N, consistent with the conclusion that locality loss is effective only when the number of experts is greater than or equal to the number of nodes. Our approach incurs slightly higher time consumption than LocMoE due to the computational overhead of token rearrangement. However, as token features converge, the required tokens gradually decrease and stabilize, leading to a decline in time consumption, which remains stable in subsequent training processes. Overall, our approach reduces training time by 2.9% to 13.3% compared to LocMoE, and by 5.4% to 46.6% compared to the Baseline.
We select 10 iterations at equal intervals from the training iterations to collect data on the time consumption of computation, communication, overlap, and idle periods, as shown in Figure 3(a). It is important to note that the data collection operation also introduces some overhead. After integrating LocMoE and our approach, the time consumption of each component decreases, with a significantly greater reduction in computation overhead compared to communication overhead. Additionally, as the cluster size increases, the proportion of computation/communication overlap decreases, and the magnitude of the reduction in computation overhead diminishes. Figure 3(b) illustrates perplexity as a measure of convergence. The convergence curves of these approaches indicate normal loss convergence, with our approach not adversely impacting convergence.

### v4

@@ Abstract
Mixture-of-Experts (MoE) architectures enable efficient scaling of large language models by activating only a subset of parameters per input. However, existing MoE models suffer from two critical limitations: (1) inefficient token-to-expert routing that causes excessive communication overhead, and (2) expert homogenization that leads to redundant computations. Current approaches address these challenges separately, failing to achieve simultaneous improvements in both training efficiency and model performance. We present Expert-Token Resonance (ETR), a theoretically-grounded bidirectional routing mechanism that fundamentally reimagines expert-token interactions in MoE architectures. Our key insight is that optimal routing requires adaptive coordination between token-choice routing (TCR) during early training phases and expert-choice routing (ECR) in later stages. We prove that this dynamic approach maximizes training success rate—the probability of correct token-expert assignments—while reducing the expert capacity lower bound by up to 40%. ETR incorporates three technical innovations: (1) an affinity-based routing architecture using Grouped Average Pooling (GrAP) that reduces computational complexity from O(d²) to O(d²/D) while maintaining orthogonality to prevent expert homogenization; (2) a bidirectional selection mechanism that enables both tokens and experts to actively participate in the routing process based on cosine similarity scores; and (3) an adaptive capacity strategy that dynamically adjusts expert bounds based on training progress, eliminating communication bubbles in All-to-All operations. Extensive experiments on Ascend NPU clusters demonstrate that ETR achieves 5.4%-46.6% improvements in end-to-end training efficiency compared to baseline MoE implementations, with 9.7%-14.5% performance gains across GDAD, GPQA, HumanEval, and TeleQnA benchmarks. These results establish ETR as the efficient MoE routing approach to deliver substantial improvements in both computational efficiency and model quality, enabling practical deployment of larger sparse models previously constrained by communication bottlenecks.
@@ Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in understanding complex semantic relationships and generating coherent text (Zhao et al. 2023). However, as model parameters scale to billions or even trillions, the computational and communication costs grow prohibitively (Jiang et al. 2024b). The Mixture-of-Experts (MoE) architecture offers a promising solution by activating only a subset of model parameters for each input, enabling efficient model scaling without proportional increases in computational demands (Lepikhin et al. 2020; Fedus, Zoph, and Shazeer 2022). Recent MoE-based LLMs, such as DeepSeek-V3 (Liu et al. 2024a) and Mixtral (Jiang et al. 2024a), have achieved state-of-the-art performance across various benchmarks.
Despite these successes, MoE models face two fundamental challenges that limit their practical deployment and effectiveness. First, the routing efficiency problem: current routing mechanisms suffer from suboptimal token-to-expert assignments, leading to significant computational waste through excessive padding in All-to-All communications and underutilized expert capacity. Second, the expert homogenization problem: existing routing strategies often fail to maintain expert specialization, resulting in redundant computations where multiple experts learn similar representations. These challenges are particularly pronounced during early training phases when routing networks lack sophisticated assignment capabilities.
Previous approaches have attempted to address these issues separately. Load balancing techniques (Zhou et al. 2022b; Dai et al. 2022) focus on distributing tokens evenly across experts but often sacrifice routing quality. Expert specialization methods (Xie et al. 2024; Li et al. 2024) aim to maintain diversity but typically incur additional computational overhead. Critically, no existing work has demonstrated how to simultaneously improve both training efficiency and model performance through a unified routing mechanism.
This paper introduces Expert-Token Resonance (ETR), a novel bidirectional routing strategy that fundamentally rethinks how tokens and experts interact in MoE architectures. Our key insight is that optimal routing requires not only tokens selecting appropriate experts (token-choice routing, TCR) but also experts actively selecting their most relevant tokens (expert-choice routing, ECR). We theoretically prove that TCR achieves higher training success rates during early training when routing networks need refinement, while ECR excels in later stages when expert specialization becomes critical. By adaptively combining both paradigms, ETR achieves what previous methods could not: simultaneous improvements in training efficiency and model performance.
Specifically, our contributions are threefold:
@@
    A theoretically-grounded bidirectional routing mechanism that dynamically balances token-choice and expert-choice routing based on training progress. We prove that this approach maximizes the training success rate—the probability of correct token-expert assignments—throughout the entire training process while reducing the required expert capacity lower bound by up to 40% compared to conventional methods.
@@
    An efficient affinity-based routing architecture leveraging Grouped Average Pooling (GrAP) layers that reduces computational complexity from $O(d^{2})$ to $O(d^{2}/D)$ compared to traditional MLP routers, where $d$ is the hidden dimension and $D$ is the grouping factor. The orthogonality properties of GrAP naturally prevent expert homogenization while enabling precise affinity computations through cosine similarity.
@@
    Comprehensive empirical validation on Ascend NPU clusters demonstrating that ETR improves end-to-end training efficiency by 5.4%-46.6% compared to baseline MoE implementations and 2.9%-13.3% compared to state-of-the-art LocMoE, while simultaneously enhancing downstream task performance by 9.7%-14.5% and 1.7%-4.1%, respectively.
The significance of this work extends beyond incremental improvements. By addressing the fundamental trade-off between routing quality and computational efficiency, ETR enables practical deployment of larger MoE models that were previously constrained by communication bottlenecks. Our theoretical analysis provides new insights into the dynamics of expert-token interactions, opening avenues for future research in sparse model architectures.
Related Work
Routing Mechanisms
Routing design fundamentally determines MoE performance. Token-choice routing (TCR) (Shazeer et al. 2017a) allows tokens to select experts but suffers from load imbalance. Expert-choice routing (ECR) (Zhou et al. 2022a) ensures balanced loads by having experts select tokens, risking the loss of critical tokens. Recent dynamic routing strategies (Huang et al. 2024) adapt expert allocation to input complexity but remain unidirectional—either tokens choose experts or vice versa. This fundamental limitation prevents simultaneous optimization of routing quality and load balance.
Expert Specialization
Expert homogenization critically limits MoE performance. Prior work addressed this through fine-grained expert segmentation (Dai et al. 2024), competitive mechanisms (Pham et al. 2024), and theoretical analyses linking data clustering to specialization (Chen et al. 2022). However, these approaches require auxiliary losses or complex training procedures.
System Optimization
System-level MoE optimizations focus on hardware utilization through block-sparse operations (Gale et al. 2022) and communication reduction via expert sharding (Balmau et al. 2025) or sequence migration (Chen et al. 2024). While these approaches improve specific bottlenecks, they overlook communication bubbles in All-to-All operations—a critical inefficiency.
Positioning of This Work
Recent surveys (Cai et al. 2024) reveal that existing MoE methods typically prioritize either routing quality or computational efficiency. Our ETR framework addresses previously overlooked communication inefficiencies and training bottleneck, suggesting potential for more efficient MoE deployment at scale.
@@
As shown in Figure 1, GrAP performs average pooling with group segmented by the number of experts. With GrAP as the layer of feature extraction, the formulation of $\bm{W}_{\text{aff}}$ is as followed:
@@
We consider our affinity-driven active selection routing as a hybrid of TCR (Clark et al. 2022; Zhou et al. 2022b) and ECR. As the name suggested, TCR lets each token choose its top-scored experts, and ECR lets each expert choose its top-scored tokens. Specifically, we use the result of the expert-token affinity metrics as the affinity score between tokens and experts. In conventional TCR routing strategy, the tokens are simply route to their Top-1 expert. In our hybird TCR+ECR routing strategy, experts also select tokens for processing from assigned tokens according to affinity scores:
@@
and then the expert to choose its Top-$\ell$ tokens where $\ell$ is determined by a threshold of the sum of affinity scores:
@@
Feed-forward network (FFN) layers are commonly employed in expert networks, allowing each expert to learn independently as a separate neural network, thus preventing interference between samples. This mechanism leads to a severe load imbalance, as experts frequently selected in the early stages are more likely to be chosen in later stages. To mitigate this skewness in token allocation, the auxiliary loss (Shazeer et al. 2017b) has been proposed. Building upon the auxiliary loss, our work introduces a loss bias term based on data locality, represented as $L_{\mathrm{loc}}=\mu\mathrm{KL}(D_{\mathrm{c}}||D_{\mathrm{l}})=-\mu\int{D_{\mathrm{c}}(x)\ln[\frac{D_{\mathrm{l}}(x)}{D_{\mathrm{c}}(x)}]\mathrm{d}x}$, i.e., the Kullback-Leibler (KL) divergence of the current distribution $D_{\mathrm{c}}(x)$ and the fully localized distribution $D_{\mathrm{l}}(x)$. This loss term serves as a soft constraint, encouraging tokens to be sent to experts residing on the same node, thereby mitigating the substantial overhead incurred by partial inter-node communication.
@@
The proposed hybrid TCR+ECR bidirectional routing mechanism operates through a two-stage process: tokens are initially assigned to experts based on the similarity between their feature fragments and the corresponding gating weights, followed by each expert selecting the Top-$\ell$ most relevant tokens through a scoring mechanism. This hybrid approach dynamically determines the number of tokens processed by each expert via an adaptive threshold, thereby ensuring the retention of class-discriminative tokens while optimizing computational efficiency.
Dynamic Lower Bound Module for Expert Capacity in ETR
To explain the motivation of our method, we show some theoretical insights in this section.
@@
Each input $\bm{x}\in\mathcal{R}^{sd}$ with $s$ tokens is comprised of one class-discriminative pattern $\bm{o}_{1},\dots,\bm{o}_{n}\in\mathcal{R}^{d}$, with each decides the label in $[n]$, and s1s-1 class-irrelevant patterns $\bm{r}\sim\mathcal{N}$ for certain distribution $\mathcal{N}$. For example, $\bm{x}=(\bm{r}_{1},\bm{r}_{2},\bm{o}_{1},\bm{r}_{3},\dots,\bm{r}_{s-1})$ has label $1$, where $\bm{r}_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N},\forall i\in[s-1]$.
@@
Phase 1: Router training This process ensures that each expert only receives the class-discriminative tokens related to the specific class.
Phase 2: Expert training This process is designed to establish each expert’s ability to handle and solve problems.
To quantitatively measure the difference between TCR and ECR, we define training success rate of input motivated by the training process of MoE.
@@
$i\sim\mathrm{Unif}([s]),\bm{x}_{i}\sim\mathrm{Unif}\left(\{\bm{o}_{1},\dots,\bm{o}_{n}\}\right).$ (7)
@@
$\mathcal{P}(\delta_{\bm{o}_{i},i}\geq\delta_{\bm{x}_{j},i},\forall j\in[s])=p_{i}\geq 1/n,\forall i\in[n].$ (8)
@@
$\mathcal{P}(\text{ECR succeed})\begin{cases}\leq\frac{1}{n}\sum_{i=1}^{n}e^{-\frac{(s-1)q_{i}}{8}},&C\leq(s-1)q_{i}/2,\\ \geq 1-e^{-3C/16},&C\geq 2sq_{i}.\end{cases}$ (12)
@@
Under the assumptions of orthogonal gating weights and uniform data distribution, we establish that the lower bound of expert capacity is given by $C_{\mathrm{min}}=\frac{1}{n}\exp\{d\delta_{\mathrm{max}}^{2}/(2-\delta_{\mathrm{max}}^{2})\}$, where the capacity is intrinsically linked to the angular relationship between gating weights and tokens.
Building on Theorem 5 and the evolution of feature distributions during training, we demonstrate the optimality of transitioning from TCR to ECR. In early training stages, when class-irrelevant tokens exhibit near-isotropic distribution with $q_{i}=\Theta(1)$, TCR achieves superior training success rates of $C/s$ compared to ECR’s exponentially decaying rate of $e^{-s}$, necessitating a larger expert capacity of $C=\Theta(s)$. As training progresses and experts develop discriminative capabilities, the distribution shifts such that $q_{i}\ll 1$ or $sq_{i}\leq C^{*}$ for some constant $C^{*}>0$, at which point ECR approaches unit success rate while TCR remains bounded by $C/s$, enabling efficient operation with reduced capacity $C=\Theta(1)$ when $C\geq 2C^{*}$.
@@
To mitigate the serial execution bottlenecks inherent in LLM training, we implement Communication Over Computation (CoC) optimization, which transforms sequentially dependent matrix multiplication and collective communication operations in parallel linear layers into unified, fine-grained kernels. By leveraging MTE’s remote memory access capabilities, CoC fuses computation and communication primitives, enabling pipeline-parallel execution that overlaps previously serialized operations.
@@
We implement our approach using Mixtral 8×7B, a 46.7B-parameter model with Group Query Attention (GQA) and 32 sparse MoE blocks. Each block contains 8 experts, with tokens routed to their top-2 selections. To accommodate long-context applications, we extend the sequence length to 32,768 tokens. Our experiments span three cluster scales with tailored parallelization strategies: 32 NPUs (TP=4, PP=4, DP=2, EP=2), 64 NPUs (TP=8, PP=4, DP=2, EP=2), and 256 NPUs (TP=8, PP=8, DP=4, EP=2), maintaining a global batch size of 128 throughout. Additional experimental details are provided in the Appendix.
@@
We consistently employ Top-1 routing to align implementation with our theoretical framework. The baseline model uses constrained expert capacity rather than groupedGEMM, preventing token dropout with a capacity factor of 1.1. LocMoE incorporates distributional uniformity and estimates expert capacity via a theoretically-derived lower bound formula from the initial batch, maintaining this value throughout training. Our approach (”LocMoE+” in figures) constrains score sum ranges, processes hidden states, and dynamically calculates expert capacity. The subsequent analysis examines training efficiency, convergence, and memory utilization across multiple Ascend cluster configurations.
@@
Figure 4 presents the training overhead analysis across the initial 1000 training iterations. To ensure measurement stability and exclude initialization artifacts, we commence time profiling from the fifth iteration onward. The baseline model exhibits consistent temporal performance throughout the evaluation period. In contrast, LocMoE demonstrates a marginal decrease in execution time as training progresses, with this trend being particularly pronounced in the 32N and 64N configurations. This observation corroborates our hypothesis that locality-aware optimization achieves optimal efficiency when the number of experts meets or exceeds the number of computational nodes. Our proposed method introduces a modest computational overhead relative to LocMoE, attributable to the token rearrangement mechanism. However, this overhead diminishes progressively as token representations converge during training. Specifically, the convergence of token features leads to a reduction in the number of tokens requiring rearrangement, resulting in stabilized computational costs in later training phases. Empirically, our approach achieves a reduction in total training time ranging from 2.9% to 13.3% compared to LocMoE, and from 5.4% to 46.6% relative to the baseline configuration.
Figure 5 depicts the temporal distribution of computational phases across training. We sample performance metrics at ten equidistant intervals throughout training, capturing computation, communication, overlap, and idle time. While profiling introduces minor overhead, this methodology provides robust insights into system behavior.
Both LocMoE and our proposed method demonstrate reduced latency across all components, with computational overhead exhibiting more substantial improvements than communication costs. This efficiency gain follows a clear pattern: as cluster size expands, the computation-communication overlap ratio decreases, accompanied by diminishing returns in computational speedup. This trend reflects the inherent scalability challenges in distributed MoE architectures.
Figure 5: The average composition of computation, communication, overlap, and idle with different schemes and cluster sizes.
Figure 6 validates that these efficiency improvements preserve model quality. All methods exhibit comparable convergence trajectories, confirming that our optimization strategy maintains training stability while delivering performance gains. The perplexity curves demonstrate that accelerated training does not compromise the fundamental learning dynamics of the model.
Figure 6: The perplexity during training iterations with different schemes.
Figure 7 presents the operator-level computational profiling across different hardware components. The system leverages AI CORE for matrix multiplication and convolution operations, AI VECTOR CORE for parallelized vector computations, MIX AIC for heterogeneous operator fusion, and AI CPU for specialized AI instruction execution. Our token selection strategy yields substantial performance gains: the FFN MatMul operator achieves a 17× speedup compared to the baseline and 2.6× improvement over LocMoE. This optimization translates to a 2.8× reduction in cumulative MatMul execution time and a 2.6× decrease in Cube computational load. The rearrangement-associated operators (TopK and IndexPutV2) exhibit marginal overhead increases, representing an acceptable trade-off for the significant computational savings achieved through selective token processing.
Figure 7: The distribution of time consumption for operators.
Figure 8 analyzes memory consumption patterns during stable training phases, based on 100,000 memory profiling samples per device. Our approach demonstrates substantial memory efficiency gains, achieving 4.57-16.27% reduction compared to the baseline and 2.86-10.5% reduction compared to LocMoE. The memory optimization exhibits scale-dependent characteristics: larger clusters show reduced computational overhead proportions and correspondingly narrower memory usage differentials. Furthermore, our method effectively eliminates transient memory spikes and reduces short-term memory fluctuations, contributing to more predictable and stable resource utilization throughout training.